please help me solve this problem:
x+3y=11
4x+y=6
using (3,-6)
may you determine if the given ordered pair is a solution to the linear equations and explain why. please and thank you
Answer by mananth(16949) 
 1 x + 3 y = 11 .............1
4 x + 1 y = 6 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by -3
1 x 3 y = 11
-12 x -3 y = -18
Add the two equations
-11 x = -7
/ -11
x = 1
plug value of x in (1)
1 x + 3 y = 11
1 + 3 y = 11
3 y = 11 -1
3 y = 10
y = 3
x= 1
y= 3
(1,3) is the solution.
If (3,-6) is a solution then it must satisfiy both equations.
plug (3,-6) in one of the equations
x+3y=11
3+(-6)3=-15.
This is not equal to 11 hence (3,-6) is not the solution of this system
m.ananth@hotmail.ca
|
|
|